Rate of heat leaving +rubric for answering question!

A window pane that measures 20.0 cm by 15.0 cm is set into the front door of a house. The temperature outdoors is -15.0° C and inside is 22.0° C. If the glass is 0.32 cm thick, at what rate does heat leave the house through the window? If the single paned window is replaced with a double paned window with an air gap of 0.50 cm between the two glass panes, each 0.32 cm thick, at what rate does heat leave the house through the double paned window? a) b)
Fails to Meet Criteria Fair Read and Translate the 2 Points 1 Points A clear re-statement of the problem is given in a list/ table of A clear partial An unclear Statement re-statement of the problem is given ina re-statement of the relevant known and unknown variables. listi table of relevant known and unknown problem is given in a list/ table of relevant known and unknown State Applicable Laws, Assumptions 3 Points 2 Points 1 Points simplifications are given and related to how they allow the concepts/laws to be used to solve this particular situation. are given with no information about how they relate to the concepts/laws that will be used, or an conceptslaws are listed with no assumptions or simplifications, or Simplifications Physical Representation A cleary labeled. 2 Points A reasonable 1 Points An incorrect correct physical representation is given, with all quantities and symbols clearly defined. given, but is not given, or one that is correct, but does not include any abels or labeled, does not define al quantities, or a clear given but it contains a Mathematical Representation 3 Points A correct mathematical representation is given in symbols with 2 Points A partially correct representation is 1 Points An incorrect representation is given in symbols with no numbers. given in symbols with no numbers. Work through Mathematics 4 Points 2 Points 1 Points
Answer #1

(a) At what rate does heat leave the house through the window?

using a formula, we have

(Q / t) = \kappa A (Tin - Tout) / d

where, \kappa = thermal conductivity for the glass = 0.75 W/m 0C

Tin = inside temperature = 22 0C

Tout = outside temperature = - 15 0C

d = thickness of the glass = 0.32 cm = 0.0032 m

A = area of the glass = (0.2 m) (0.15 m) = 0.03 m2

then, we get

(Q / t) = (0.75 W/m 0C) (0.03 m2) [(22 0C) + (15 0C)] / (0.0032 m)

(Q / t) = [(0.8325 W.m) / (0.0032 m)]

(Q / t) = 260.1 W